The perimeter of the rectangular playing field is 430 yards. The length of the field is 5 yards less than triple the width. What are the dimensions of the playing field?
Answer by addingup(3677)

2L+2W = 430
L = 3W-5
2(3W-5)+2W = 430
6W-10+2W = 430
8W = 440
W = 55 and L = 3(55)-5 = 160
----------------------------
Check:
2L+2W = 430
2(160)+2(55) = 430
320+110 = 430 Correct.
:
John